# USTC-ICTS/PCFT-25-37

### v1

@@ Appendix A The topsector of sunset
$\left.+2\eta_{1}{\partial\over\partial\eta_{3}}{\partial\over\partial\eta_{1}}G_{111}\right\}+$$penalty 10000\ \penalty 10000\$
A crucial point for the group II and III is that when we consider the massless limit, some degree two relations will be reduced to the degree one relations, such as (33) and (35). This difference will lead to different master integrals, which have been discussed in the main text.
To finish the reduction of the set ${\cal U}_{1}$ for massive case, we need to generate one more equation, which comes from ${\partial\over\partial\eta_{1}}({\text{(I-2)+(I-3)}})$ and is given by
${2m_{1}^{2}(K^{2}-m_{1}^{2})\over s_{0}^{2}}{\partial^{2}\over\partial\eta_{1}^{2}}G_{111}$ (37)
$+{m_{2}^{2}(f_{-++}-2(m_{1}^{2}-m_{3}^{2}))\over s_{0}^{2}}{\partial\over\partial\eta_{2}}{\partial\over\partial\eta_{1}}G_{111}$
$+{m_{3}^{2}(f_{-++}-2(m_{1}^{2}-m_{2}^{2}))\over s_{0}^{2}}{\partial\over\partial\eta_{3}}{\partial\over\partial\eta_{1}}G_{111}$
$=$ $\left\{{-3(D-3)m_{3}^{2}\over s_{0}}{\partial\over\partial\eta_{3}}G_{111}+{-3(D-3)m_{2}^{2}\over s_{0}}{\partial\over\partial\eta_{2}}G_{111}\right$.
$-{2(K^{2}-2m_{1}^{2})\over s_{0}}\eta_{1}{\partial^{2}\over\partial\eta_{1}^{2}}G_{111}$
$+\left({(D-3)(m_{2}^{2}+m_{3}^{2})+(D-4)K^{2}\over s_{0}}\right$.
$\left.+{(16-5D)m_{1}^{2}\over s_{0}}\right){\partial\over\partial\eta_{1}}G_{111}$
$+\left({(-f_{-+-}+2(m_{1}^{2}-m_{2}^{2}))\eta_{3}\over s_{0}}\right$.
$\left.+{3m_{3}^{2}(\eta_{1}-\eta_{2}-\eta_{3})\over s_{0}}\right){\partial\over\partial\eta_{3}}{\partial\over\partial\eta_{1}}G_{111}$
$+\left({(-f_{--+}+2(m_{1}^{2}-m_{3}^{2}))\eta_{2}\over s_{0}}\right$.
$\left.\left.+{3m_{2}^{2}(\eta_{1}-\eta_{2}-\eta_{3})\over s_{0}}\right){\partial\over\partial\eta_{2}}{\partial\over\partial\eta_{1}}G_{111}\right\}+.$...
where again only nonzero contribution terms for reducing the set $(n_{1},n_{2},n_{3},0,0)$ have been kept.
Appendix B The top sector of massless double box
From $10$ fundamental IBP relations, we can solve the following $4$ degree one operators
${s\over s_{0}}{\partial\over\partial\eta_{3}}G=-\left\{(D-5)G-2\eta_{1}{\partial\over\partial\eta_{1}}G-(\eta_{1}+\eta_{2}){\partial\over\partial\eta_{2}}G\right$.
$\left.-(\eta_{1}+\eta_{3}){\partial\over\partial\eta_{3}}G+(\eta_{4}-\eta_{1}-\eta_{5}){\partial\over\partial\eta_{5}}G+\eta_{8}{\partial\over\partial\eta_{8}}G\right\}\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\$(38)
${s\over s_{0}}{\partial\over\partial\eta_{6}}G=-\left\{(D-5)G+(\eta_{1}-\eta_{4}-\eta_{5}){\partial\over\partial\eta_{5}}G+\eta_{9}{\partial\over\partial\eta_{9}}G\right$.
$\left.-2\eta_{4}{\partial\over\partial\eta_{4}}G-(\eta_{4}+\eta_{6}){\partial\over\partial\eta_{6}}G-(\eta_{4}+\eta_{7}){\partial\over\partial\eta_{7}}G\right\}\penalty 10000\ \penalty 10000\$(39)
${s\over s_{0}}{\partial\over\partial\eta_{1}}G=-\left\{(D-5)G-(\eta_{1}+\eta_{3}){\partial\over\partial\eta_{1}}G-(\eta_{2}+\eta_{3}){\partial\over\partial\eta_{2}}G\right$.
$\left.-2\eta_{3}{\partial\over\partial\eta_{3}}G+(\eta_{6}-\eta_{3}-\eta_{5}){\partial\over\partial\eta_{5}}G+\eta_{8}{\partial\over\partial\eta_{8}}G\right\}$
$+\left\{{s\over 2s_{0}}\eta_{8}G\right\}\penalty 10000\$(40)
${s\over s_{0}}{\partial\over\partial\eta_{4}}G=\left\{-(D-5)G+(\eta_{6}+\eta_{4}){\partial\over\partial\eta_{4}}G\right$.
$\left.+(\eta_{6}-\eta_{3}+\eta_{5}){\partial\over\partial\eta_{5}}G+(\eta_{6}+\eta_{7}){\partial\over\partial\eta_{7}}G\right$.
$\left.+2\eta_{6}{\partial\over\partial\eta_{6}}G-\eta_{9}{\partial\over\partial\eta_{9}}G\right\}+\left\{{-s\over 2s_{0}}\eta_{9}G\right\}\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\$(41)
and $6$ degree two operators
$2{\partial\over\partial\eta_{9}}{\partial\over\partial\eta_{5}}G=-\left\{(D-5)G-(\eta_{1}+\eta_{2}){\partial\over\partial\eta_{1}}G\right$.
$\left.-2\eta_{2}{\partial\over\partial\eta_{2}}G-(\eta_{2}+\eta_{3}){\partial\over\partial\eta_{3}}G+(\eta_{4}-\eta_{2}-\eta_{5}){\partial\over\partial\eta_{5}}G\right$.
$\left.+\eta_{8}{\partial\over\partial\eta_{8}}G\right\}+\left\{{(s+t)\over 2s_{0}}\eta_{8}G\right\}\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\$(42)
$2{\partial\over\partial\eta_{8}}{\partial\over\partial\eta_{5}}G=\left\{{s\over s_{0}}{\partial\over\partial\eta_{5}}G\right\}+\left\{-(D-5)G\right$.
$\left.+(\eta_{7}+\eta_{4}){\partial\over\partial\eta_{4}}G+(\eta_{7}-\eta_{3}+\eta_{5}){\partial\over\partial\eta_{5}}G+2\eta_{7}{\partial\over\partial\eta_{7}}G\right$.
$\left.+(\eta_{7}+\eta_{6}){\partial\over\partial\eta_{6}}G-\eta_{9}{\partial\over\partial\eta_{9}}G\right\}+\left\{{t\over 2s_{0}}\eta_{9}G\right\}\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\$(43)
$\left\{2{\partial\over\partial\eta_{9}}{\partial\over\partial\eta_{2}}G\right\}=\left\{-(D-5)G+(\eta_{5}+\eta_{1}-\eta_{4}){\partial\over\partial\eta_{1}}G\right$.
$\left.+(\eta_{5}+\eta_{2}-\eta_{4}){\partial\over\partial\eta_{2}}G+(\eta_{5}+\eta_{3}-\eta_{6}){\partial\over\partial\eta_{3}}G\right$.
$\left.+2\eta_{5}{\partial\over\partial\eta_{5}}G-\eta_{8}{\partial\over\partial\eta_{8}}G\right\}+\left\{{s\over 2s_{0}}\eta_{8}G\right\}$
$+\left\{{1\over 2}(\eta_{7}-\eta_{6})\eta_{8}G\right\}\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\$(44)
$\left\{2{\partial\over\partial\eta_{8}}{\partial\over\partial\eta_{7}}G\right\}=\left\{{s\over s_{0}}{\partial\over\partial\eta_{7}}G\right\}+\left\{-(D-5)G+2\eta_{5}{\partial\over\partial\eta_{5}}G\right$.
$\left.+(\eta_{5}-\eta_{1}+\eta_{4}){\partial\over\partial\eta_{4}}G+(\eta_{5}-\eta_{3}+\eta_{6}){\partial\over\partial\eta_{6}}G\right$.
$\left.+(\eta_{5}-\eta_{3}+\eta_{7}){\partial\over\partial\eta_{7}}G-\eta_{9}{\partial\over\partial\eta_{9}}G\right\}$
$+\left\{{1\over 2}(\eta_{2}-\eta_{1})\eta_{9}G\right\}\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\$(45)
$2{\partial\over\partial\eta_{8}}{\partial\over\partial\eta_{2}}G=\left\{{s+t\over s_{0}}{\partial\over\partial\eta_{2}}G+{s\over s_{0}}{\partial\over\partial\eta_{3}}G\right$.
$\left.+{4(D-4)s_{0}\over s}{\partial\over\partial\eta_{8}}G+{4s_{0}\over s}\eta_{8}{\partial^{2}\over\partial\eta_{8}^{2}}G\right$.
$\left.-{2s_{0}\over s}(\eta_{1}+\eta_{3}-\eta_{4}+2\eta_{5}-\eta_{6}){\partial\over\partial\eta_{8}}{\partial\over\partial\eta_{5}}G\right$.
$\left.-{2s_{0}\over s}(\eta_{1}+\eta_{3}){\partial\over\partial\eta_{8}}{\partial\over\partial\eta_{3}}G\right$.
$\left.-{2s_{0}\over s}(\eta_{1}+2\eta_{2}+\eta_{3}){\partial\over\partial\eta_{8}}{\partial\over\partial\eta_{2}}G\right$.
$\left.-{2s_{0}\over s}(3\eta_{1}+\eta_{3}){\partial\over\partial\eta_{8}}{\partial\over\partial\eta_{1}}G\right\}+\left\{(D-6)G\right$.
$\left.-(\eta_{4}+\eta_{7}){\partial\over\partial\eta_{4}}G+(\eta_{3}-\eta_{5}-\eta_{6}){\partial\over\partial\eta_{5}}G\right$.
$\left.-(\eta_{6}+\eta_{7}){\partial\over\partial\eta_{6}}G-2\eta_{7}{\partial\over\partial\eta_{7}}G\right$.
$\left.-\eta_{8}{\partial\over\partial\eta_{8}}G+\eta_{9}{\partial\over\partial\eta_{9}}G\right\}+\left\{-{t\over 2s_{0}}\eta_{9}G\right\}\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\$(46)
$2{\partial\over\partial\eta_{9}}{\partial\over\partial\eta_{7}}=\left\{-{s\over s_{0}}{\partial\over\partial\eta_{6}}G+{t\over s_{0}}{\partial\over\partial\eta_{7}}G+{4s_{0}\over s}\eta_{9}{\partial^{2}\over\partial\eta_{9}^{2}}G\right$.
$\left.+{4(D-4)s_{0}\over s}{\partial\over\partial\eta_{9}}G-{2s_{0}\over s}(3\eta_{4}+\eta_{6}){\partial\over\partial\eta_{9}}{\partial\over\partial\eta_{4}}G\right$.
$\left.+{2s_{0}\over s}(\eta_{1}+\eta_{3}-\eta_{4}-2\eta_{5}-\eta_{6}){\partial\over\partial\eta_{9}}{\partial\over\partial\eta_{5}}G\right$.
$\left.-{2s_{0}\over s}(\eta_{4}+3\eta_{6}){\partial\over\partial\eta_{9}}{\partial\over\partial\eta_{6}}G\right$.
$\left.-{2s_{0}\over s}(\eta_{4}+\eta_{6}+2\eta_{7}){\partial\over\partial\eta_{9}}{\partial\over\partial\eta_{7}}G\right\}$
$+\left\{(D-4)G-(\eta_{1}+\eta_{2}){\partial\over\partial\eta_{1}}G-2\eta_{2}{\partial\over\partial\eta_{2}}G+\eta_{9}{\partial\over\partial\eta_{9}}G\right$.
$\left.-(\eta_{2}+\eta_{3}){\partial\over\partial\eta_{3}}G-(\eta_{1}-\eta_{4}+\eta_{5}){\partial\over\partial\eta_{5}}G+\eta_{8}{\partial\over\partial\eta_{8}}G\right\}$
$+\left\{-{(s+t)\over 2s_{0}}\eta_{8}G\right\}\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\$(47)
where for simplicity we have neglected the contributions from subsectors.
Now we construct new equations according to step (B) (or more directly, just using the integrability conditions to keep equations up to degree two). After simplification using known reduction rules, when applying to the set ${\cal U}_{2}=(0,0,0,0,0,0,0,n_{8},n_{9})$, the first group, i.e., ${\partial\over\partial\eta_{9}}\left({\partial\over\partial\eta_{8}}{\partial\over\partial\eta_{i}}\right)-{\partial\over\partial\eta_{8}}\left({\partial\over\partial\eta_{9}}{\partial\over\partial\eta_{i}}\right)=0,i=2,5,7$ gives
$0$ $=$ $\left\{{2(D-4)s_{0}\over s}{\partial\over\partial\eta_{8}}{\partial\over\partial\eta_{9}}G+{2s_{0}\over s}\eta_{8}{\partial^{2}\over\partial\eta_{8}^{2}}{\partial\over\partial\eta_{9}}G\right\}$ (48)
$\left\{+{(D-4)\over 2}{\partial\over\partial\eta_{8}}G+{1\over 2}\eta_{8}{\partial^{2}\over\partial\eta_{8}^{2}}G+{1\over 2}\eta_{9}{\partial^{2}\over\partial\eta_{9}^{2}}G\right$.
$\left.-\eta_{8}{\partial\over\partial\eta_{8}}{\partial\over\partial\eta_{9}}G\right\}+R_{1}\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\$
$0$ $=$ $\left\{+{(D-4)\over 2}{\partial\over\partial\eta_{8}}G+{1\over 2}\eta_{8}{\partial^{2}\over\partial\eta_{8}^{2}}G-{(D-4)\over 2}{\partial\over\partial\eta_{9}}G\right$. (49)
$\left.-{1\over 2}\eta_{9}{\partial^{2}\over\partial\eta_{9}^{2}}G\right\}+R_{2}\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\$
$0$ $=$ $\left\{{-2(D-4)s_{0}\over s}{\partial\over\partial\eta_{8}}{\partial\over\partial\eta_{9}}G-{2s_{0}\over s}\eta_{9}{\partial^{2}\over\partial\eta_{8}}{\partial^{2}\over\partial\eta_{9}^{2}}G\right\}$ (50)
$\left\{-{(D-4)}{\partial\over\partial\eta_{8}}G-{1\over 2}\eta_{8}{\partial^{2}\over\partial\eta_{8}^{2}}G\right$.
$\left.+{(D-4)\over 2}{\partial\over\partial\eta_{9}}G+{1\over 2}\eta_{9}{\partial^{2}\over\partial\eta_{9}^{2}}G\right$.
$\left.+\left(-\eta_{9}-{4(D-4)s_{0}^{2}\over s^{2}}(\eta_{4}+\eta_{6})\right){\partial\over\partial\eta_{8}}{\partial\over\partial\eta_{9}}G\right$.
$\left.-{4s_{0}^{2}\over s^{2}}(\eta_{4}+\eta_{6})\eta_{9}{\partial\over\partial\eta_{8}}{\partial^{2}\over\partial\eta_{9}^{2}}G\right\}+R_{3}\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\$
where $R_{i}$ are terms with degree zero or lower. When writing these three equations into the recurrence form, we have
$-R_{1}$ $=$ ${2s_{0}\over s}(n_{8}+1)(n_{9}+1)(D-4+n_{8})F_{\vec{n}+\vec{e}_{8}+\vec{e}_{9}}$ (51)
$+{1\over 2}(n_{8}+1)(D-4+n_{8})F_{\vec{n}+\vec{e}_{8}}$
$+{1\over 2}(n_{9}+1)(n_{9}-2n_{8})F_{\vec{n}+\vec{e}_{9}}\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\$
$-R_{2}$ $=$ ${1\over 2}(n_{8}+1)(D-4+n_{8})F_{\vec{n}+\vec{e}_{8}}$ (52)
$-{1\over 2}(n_{9}+1)(D-4+n_{9})F_{\vec{n}+\vec{e}_{9}}\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\$
$R_{3}$ $=$ ${2s_{0}\over s}(n_{8}+1)(n_{9}+1)(D-4+n_{9})F_{\vec{n}+\vec{e}_{8}+\vec{e}_{9}}$ (53)
$+{1\over 2}(n_{8}+1)(2(D-4)+n_{8}+2n_{9})F_{\vec{n}+\vec{e}_{8}}$
$-{1\over 2}(n_{9}+1)(D-4+n_{9})F_{\vec{n}+\vec{e}_{9}}\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\$
We can solve $F_{\vec{n}+\vec{e}_{8}+\vec{e}_{9}},F_{\vec{n}+\vec{e}_{8}},F_{\vec{n}+\vec{e}_{9}}$ as long as above three relations are linear independent. This can be checked since the coefficient determinant is not zero for any values of $n_{8},n_{9}$.
Equations from the second and third groups are
$0$ $=$ $\left\{{t\over 2s_{0}}{\partial\over\partial\eta_{2}}{\partial\over\partial\eta_{5}}G\right\}+\left\{+{(D-5)\over 2}{\partial\over\partial\eta_{2}}G-\eta_{7}{\partial\over\partial\eta_{7}}{\partial\over\partial\eta_{2}}G\right$. (54)
$\left.+(D-{11\over 2}){\partial\over\partial\eta_{5}}G-\eta_{7}{\partial\over\partial\eta_{7}}{\partial\over\partial\eta_{5}}G-\eta_{5}{\partial^{2}\over\partial\eta_{5}^{2}}G\right$.
$\left.-{t(\eta_{1}+2\eta_{2}+\eta_{3})+s(\eta_{2}+\eta_{5}+\eta_{7})\over 2s}{\partial\over\partial\eta_{2}}{\partial\over\partial\eta_{5}}G\right\}$
$+R_{4}\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\$
$0$ $=$ $\left\{{t\over 2s_{0}}{\partial\over\partial\eta_{2}}{\partial\over\partial\eta_{7}}G\right\}+\left\{+{(D-5)\over 2}{\partial\over\partial\eta_{2}}G-\eta_{5}{\partial\over\partial\eta_{5}}{\partial\over\partial\eta_{2}}G\right$. (55)
$\left.+(D-{11\over 2}){\partial\over\partial\eta_{7}}G-\eta_{5}{\partial\over\partial\eta_{7}}{\partial\over\partial\eta_{5}}G-\eta_{7}{\partial^{2}\over\partial\eta_{7}^{2}}G\right$.
$\left.-{1\over 2}(\eta_{2}+\eta_{5}+\eta_{7}){\partial\over\partial\eta_{2}}{\partial\over\partial\eta_{7}}G\right\}+R_{5}\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\$
$0$ $=$ $\left\{+{(D-6)\over 2}{\partial\over\partial\eta_{5}}G-\eta_{5}{\partial^{2}\over\partial\eta_{5}^{2}}G-{(D-6)\over 2}{\partial\over\partial\eta_{7}}G\right$. (56)
$\left.+\eta_{7}{\partial^{2}\over\partial\eta_{7}^{2}}G\right\}+R_{6}\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\$
$0$ $=$ $\left\{+{(D-6)\over 2}{\partial\over\partial\eta_{2}}G-\eta_{2}{\partial^{2}\over\partial\eta_{2}^{2}}G-{(D-6)\over 2}{\partial\over\partial\eta_{5}}G\right$. (57)
$\left.+\eta_{5}{\partial^{2}\over\partial\eta_{5}^{2}}G\right\}+R_{7}\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\$
$0$ $=$ $\left\{-{t\over 2s_{0}}{\partial\over\partial\eta_{2}}{\partial\over\partial\eta_{7}}G\right\}+\left\{({11\over 2}-D){\partial\over\partial\eta_{2}}G+\eta_{2}{\partial^{2}\over\partial\eta_{2}^{2}}G\right$. (58)
$\left.-{(D-5)\over 2}{\partial\over\partial\eta_{7}}G+\eta_{5}{\partial\over\partial\eta_{5}}{\partial\over\partial\eta_{7}}G\right$.
$\left.+{1\over 2}(\eta_{2}+\eta_{5}+\eta_{7}){\partial\over\partial\eta_{2}}{\partial\over\partial\eta_{7}}G+\eta_{5}{\partial\over\partial\eta_{2}}{\partial\over\partial\eta_{5}}G\right\}+R_{8}\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\$
$0$ $=$ $\left\{-{t\over 2s_{0}}{\partial\over\partial\eta_{5}}{\partial\over\partial\eta_{7}}G\right\}+\left\{-(D-{11\over 2}){\partial\over\partial\eta_{5}}G+\eta_{5}{\partial^{2}\over\partial\eta_{5}^{2}}G\right$. (59)
$\left.+\eta_{2}{\partial\over\partial\eta_{2}}{\partial\over\partial\eta_{5}}G-{(D-5)\over 2}{\partial\over\partial\eta_{7}}G+\eta_{2}{\partial\over\partial\eta_{2}}{\partial\over\partial\eta_{7}}G\right$.
$\left.+{1\over 2}(\eta_{2}+\eta_{5}+\eta_{7}){\partial\over\partial\eta_{5}}{\partial\over\partial\eta_{7}}G\right\}+R_{9}\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\$
Using these $6$ relations, first we can solve three degree two operators ${\partial\over\partial\eta_{2}}{\partial\over\partial\eta_{5}}$, ${\partial\over\partial\eta_{2}}{\partial\over\partial\eta_{7}}$ and ${\partial\over\partial\eta_{5}}{\partial\over\partial\eta_{7}}$. Secondly, we have three degree one relations, which contain three degree one operators with index (0,1,0,0,0,0,0,0,0)(0,1,0,0,0,0,0,0,0),$(0,0,0,0,1,0,0,0,0)$, and $(0,0,0,0,0,0,1,0,0)$. However, they are not independent, as can be seen in the sum of (55), (56), (57), and (58), the degree two and degree one parts have been completely canceled.
Now we can use (56) to solve, for example
${(D-6)\over 2}{\partial\over\partial\eta_{7}}G-\eta_{7}{\partial^{2}\over\partial\eta_{7}^{2}}G$ (60)
$=$ ${(D-6)\over 2}{\partial\over\partial\eta_{5}}G-\eta_{5}{\partial^{2}\over\partial\eta_{5}^{2}}G+R_{6}\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\$
and
${(D-6)\over 2}{\partial\over\partial\eta_{2}}G-\eta_{2}{\partial^{2}\over\partial\eta_{2}^{2}}G$ (61)
$=$ ${(D-6)\over 2}{\partial\over\partial\eta_{5}}G-\eta_{5}{\partial^{2}\over\partial\eta_{5}^{2}}G-R_{7}\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\$
Since the left-hand side is the degree one operators, using them, we can reduce all $n_{2},n_{7}$ to zero. After iterating using these two, the unreduced lattice points are just $(0,0,0,0,n_{5},0,0,0,0)$, which contains infinity number of points.
To further reduce points $(0,0,0,0,n_{5},0,0,0,0)$, we need to produce a new relation using newly found reduction rules. Acting $\partial\over\partial\eta_{5}$} at the both sides of (61) we get
${(D-8)\over 2}{\partial^{2}\over\partial\eta_{5}^{2}}G-\eta_{5}{\partial^{3}\over\partial\eta_{5}^{3}}G$ (62)
$=$ ${(D-6)\over 2}{\partial\over\partial\eta_{5}}{\partial\over\partial\eta_{2}}G-\eta_{2}{\partial^{2}\over\partial\eta_{2}^{2}}{\partial\over\partial\eta_{5}}G+{\partial\over\partial\eta_{5}}R_{7}\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\$
Using (55), the first two terms at the right-hand side can be further reduced, and we get a perfect degree two reduction rule for $(0,0,0,0,n_{5},0,0,0,0)$. Furthermore, since the left hand side is $F_{\vec{n}+2\vec{e}_{5}}$, we can only reduce $n_{5}\geq 2$. Now the only points that can not be reduced are $n_{5}=0,1$ and we have found two master integrals.
Appendix C The top sector of Two-loop massless non-planer
Using the $10$ fundamental IBP relations, we can solve $2$ degree one operators
${s\over s_{0}}{\partial\over\partial\eta_{3}}G=\left\{-2(D-5)G-\eta_{9}{\partial\over\partial\eta_{9}}G-\eta_{8}{\partial\over\partial\eta_{8}}G\right$.
$\left.+2\eta_{7}{\partial\over\partial\eta_{7}}G+(\eta_{6}+\eta_{7}){\partial\over\partial\eta_{6}}G+(\eta_{4}+\eta_{5}){\partial\over\partial\eta_{5}}G+2\eta_{4}{\partial\over\partial\eta_{4}}G\right$.
$\left.+(\eta_{1}+\eta_{3}){\partial\over\partial\eta_{3}}G+(\eta_{1}+\eta_{2}){\partial\over\partial\eta_{2}}G+2\eta_{1}{\partial\over\partial\eta_{1}}G\right\}\penalty 10000\ \penalty 10000\ \penalty 10000\$(63)
${s\over s_{0}}{\partial\over\partial\eta_{1}}G=\left\{-2(D-5)G-\eta_{9}{\partial\over\partial\eta_{9}}G-\eta_{8}{\partial\over\partial\eta_{8}}G\right$.
$\left.+(\eta_{6}+\eta_{7}){\partial\over\partial\eta_{7}}G+2\eta_{6}{\partial\over\partial\eta_{6}}G+2\eta_{5}{\partial\over\partial\eta_{5}}G+(\eta_{4}+\eta_{5}){\partial\over\partial\eta_{4}}G\right$.
$\left.+2\eta_{3}{\partial\over\partial\eta_{3}}G+(\eta_{2}+\eta_{3}){\partial\over\partial\eta_{2}}G+(\eta_{1}+\eta_{3}){\partial\over\partial\eta_{1}}G\right\}$
$+\left\{{-s\eta_{8}-t\eta_{9}\over 2s_{0}}G\right\}\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\$(64)
and $6$ degree two operators
$2{\partial\over\partial\eta_{6}}{\partial\over\partial\eta_{8}}G=\left\{{-s\over s_{0}}{\partial\over\partial\eta_{6}}G\right\}+\left\{(D-5)G+\eta_{9}{\partial\over\partial\eta_{9}}G\right$.
$\left.+(\eta_{1}-\eta_{4}-\eta_{7}){\partial\over\partial\eta_{7}}G+(\eta_{3}-\eta_{4}-\eta_{6}){\partial\over\partial\eta_{6}}G\right$.
$\left.+(-\eta_{4}-\eta_{5}){\partial\over\partial\eta_{5}}G-2\eta_{4}{\partial\over\partial\eta_{4}}G\right\}\penalty 10000\ \penalty 10000\ \penalty 10000\$(65)
$2{\partial\over\partial\eta_{7}}{\partial\over\partial\eta_{8}}G=\left\{(D-5)G-\eta_{9}{\partial\over\partial\eta_{9}}G+2\eta_{5}{\partial\over\partial\eta_{5}}G\right$.
$\left.+(-\eta_{1}+\eta_{5}+\eta_{7}){\partial\over\partial\eta_{7}}G+(-\eta_{3}+\eta_{5}+\eta_{6}){\partial\over\partial\eta_{6}}G\right$.
$\left.+(\eta_{4}+\eta_{5}){\partial\over\partial\eta_{4}}G\right\}+\left\{-{t\over 2s_{0}}\eta_{9}G\right\}\penalty 10000\ \penalty 10000\ \penalty 10000\$(66)
$2{\partial\over\partial\eta_{5}}{\partial\over\partial\eta_{8}}G=\left\{(D-5)G-\eta_{9}{\partial\over\partial\eta_{9}}G+(2\eta_{7}){\partial\over\partial\eta_{7}}G\right$.
$\left.+(\eta_{6}+\eta_{7}){\partial\over\partial\eta_{6}}G+(-\eta_{1}+\eta_{5}+\eta_{7}){\partial\over\partial\eta_{5}}G\right$.
$\left.+(-\eta_{1}+\eta_{4}+\eta_{7}){\partial\over\partial\eta_{4}}G\right\}+\left\{{\eta_{1}-\eta_{2}\over 2}\eta_{9}G\right\}\penalty 10000\ \penalty 10000\ \penalty 10000\$(67)
[…]

### v2

@@ Appendix A The topsector of sunset
$\left(\begin{array}[]{c| c |c |c| c| l}{\partial\over\partial\eta_{1}}&{\partial\over\partial\eta_{2}}&{\partial\over\partial\eta_{3}}&{\partial\over\partial\eta_{4}}&{\partial\over\partial\eta_{3}}&\\ \hline\cr{2m_{1}^{2}(m_{2}^{2}-m_{3}^{2})\over s_{0}^{2}}&{m_{2}^{2}f_{++-}\over s_{0}^{2}}&-{m_{3}^{2}f_{+-+}\over s_{0}^{2}}&-\widehat{O}_{0;4,1}&\widehat{O}_{0;5,2}&{\rm(I-1)}\\ \hline\cr{m_{1}^{2}f_{++-}\over s_{0}^{2}}&{2m_{2}^{2}(m_{1}^{2}-m_{3}^{2})\over s_{0}^{2}}&-{m_{3}^{2}f_{-++}\over s_{0}^{2}}&\widehat{O}_{0;4,2}&-\widehat{O}_{0;5,1}&{\rm(I-2)}\\ \hline\cr{m_{1}^{2}f_{+-+}\over s_{0}^{2}}&-{m_{2}^{2}f_{-++}\over s_{0}^{2}}&{2m_{3}^{2}(m_{1}^{2}-m_{2}^{2})\over s_{0}^{2}}&-\widehat{O}_{0;4,1}&\widehat{O}_{0;5,1}&{\rm(I-3)}\\ \end{array}\right)\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\$(70)
Using (70), we can solve ${\partial\over\partial\eta_{a}},a=4,5$ by ${\partial\over\partial\eta_{i}},i=1,2,3$. Using (65), we can solve other $5$ degree-2 operators, using, for example, $\partial^{2}\over\partial\eta_{3}^{2}$}.
Checking module of second round for massive case: Using reduction rules ${\partial\over\partial\eta_{a}},a=4,5$, we can reduce $n_{4},n_{5}$ to zero. Using five degree-2 reduction rules, the points can not be reduced is the set $(0,0,n_{3},0,0)$ and points $(1,0,0,0,0)$, $(0,1,0,0,0)$. Again we need go to the third round computation.
Third round computation of massive case: Because only $n_{3}$ having not fully reduced, we just consider the action of $\partial\over\partial\eta_{3}$}. Acting it on the reduction of, for example, the reduction rule of $\partial\over\partial\eta_{4}$}, we can find the reduction rule of $\partial^{2}\over\partial\eta_{3}^{2}$}. Using this one, the un-reducible points are only following four, i.e., $(1,0,0,0,0)$, $(0,1,0,0,0)$, $(0,0,1,0,0)$ and $(0,0,0,0,0)$, which are exactly the master integrals. Thus we have found complete reduction rules.
Appendix B Example 2: The top sector of massless double box
where the last two are ISPs. Having presented the example of sunset with some details, we will be more briefly for this example.
The first round computations: The initial seed is $10$ DE’s from IBP, where $2$ are degree-1 and $8$ degree-2. The combination of degree-2 DE’s gives $2$ degree-1 DE’s. So eventually we can find $4$ T1A-type reduction rules for operators ${\partial\over\partial\eta_{i}},i=1,3,4,6$. From the remaining $6$ degree-2 DE’s, we can solve $6$ T1A-type reduction rules for ${\partial\over\partial\eta_{a}}{\partial\over\partial\eta_{j}}$, $a=8,9$ and $j=2,5,7$. Using these $10$ reduction rules, the irreducible points are $(0,n_{2},0,0,n_{5},0,n_{7},0,0)$ and $(0,0,0,0,0,0,0,n_{8},n_{9})$.
The second round computations: Since the components $n_{1},n_{3},n_{4},n_{6}$ have been fully reduced, we consider the action ${\partial\over\partial\eta_{i}},i=2,5,7,8,9$ on above $10$ reduction rules. After simplification using descendant reduction rules, we will get $9$ new non-trivial degenerated equations, which can also be understood as the result of integrability conditions.
With proper combinations, we have $4$ degree-2 DE’s and $5$ degree-1 DE’s. Using degree-2 DE’s, we solve $3$ T1A-type reduction rules for degree-2 operators $\partial^{2}\over\partial\eta_{2}\partial\eta_{5}$}, $\partial^{2}\over\partial\eta_{2}\partial\eta_{7}$} and $\partial^{2}\over\partial\eta_{5}\partial\eta_{7}$} and $1$ T1B-reduction rule for the operator $\partial^{2}\over\partial\eta_{8}\partial\eta_{9}$}.
For the $5$ degree-1 DE’s, only $3$ of them contains good operators. Using them, we can solve $3$ T2B-reduction rules, i.e., using $\partial\over\partial\eta_{8}$} to solve $\partial\over\partial\eta_{9}$} and using $\partial\over\partial\eta_{5}$} to solve $\partial\over\partial\eta_{2}$} and $\partial\over\partial\eta_{7}$}.
With solved operators $\partial\over\partial\eta_{i}$}, $i=2,7,8,9$, we can reduce $n_{2},n_{7},n_{8},n_{9}$ to zero. Finally the irreducible lattice points are just $(0,0,0,0,n_{5},0,0,0,0)$ and $(0,0,0,0,0,0,0,1,0)$.
[…]